OpenStudy (anonymous):

What is an equation of the line that passes through (0, 8) and (4, 0)?

OpenStudy (anonymous):

a.y = -2x + 4 b.y = 2x + 4 c.y = -2x + 8 d.y = 2x + 8

OpenStudy (anonymous):

Who answers correctly I will become a fan of u